OpenStudy (anonymous):

What is an equation for the line with slope 2/3 and y-intercept 9?

OpenStudy (anonymous):

y=2x/3+9 =>3y=2x+27

OpenStudy (anonymous):

y=2/3x+9

OpenStudy (anonymous):

\(y=mx+b\) you have \(m=\frac{2}{3}, b=9\)